# Tensor rank and dimension expanders

### v1

@@ 1 Overview
An $(n_{1},n_{2},n_{3})$-tensor over $\mathbb{F}$} is a function $T:[n_{1}]\times[n_{2}]\times[n_{3}]\rightarrow{\mathbb{F}}$. We can form two dimensional ‘slices’ of the tensor $T$ by fixing one of the variables. For example, we denote by T(a,,)T(a,*,*) the $n_{2}\times n_{3}$ matrix obtained by fixing $i_{1}$ to $a\in{\mathbb{F}}$. A tensor $T$ has rank one if it is of the form $T(i_{1},i_{2},i_{3})=f_{1}(i_{1})f_{2}(i_{2})f_{3}(i_{3})$ for some functions $f_{j}:[n_{j}]\rightarrow{\mathbb{F}}$. The rank of a general tensor $T$, denoted $R(T)$ is defined to be the minimal $r$ such that $T$ can be written as a sum of $r$ rank-one tensors. [^1] There are other notions of rank for tensors, all generalizing matrix rank, but the one we defined is typically called just ‘tensor rank’. When $\mathbb{F}$} is the real or complex numbers we also define the border rank of $T$, denoted $\underline{R}(T)$ to be the smallest $r$ such that there is a sequence of rank $\leq r$ tensors converging to TT. [^2] One can define border rank also over fields of positive characteristic as being in the Zariski closure of low rank tensors. While it is easy to show the existence of tensors of high rank (as high as $\Omega(n^{2})$ for $(n,n,n)$ tensors) it is a longstanding open problem to construct high-rank tensors explicitly. By explicit construction we mean that, there is a deterministic algorithm that, given $n$, runs in time $\textsf{poly}(n)$ and outputs an $(n,n,n)$-tensor $T_{n}$ of the required rank (more formally, an explicit construction produces an infinite family of tensors, not a single one). The difficulty in constructing explicit tensors of high rank stems from the fact that such constructions would lead to lower bounds on arithmetic circuits, a notoriously difficult open problem in theoretical computer science (see [SY10] for more details on this connection).
The currently best known explicit lower bounds on tensor rank are those in [AFT11] and are of the form $3n-O(\log_{2}(n))$ for an $(n,n,n)$-tensor (over any field $\mathbb{F}$}). Their work also includes a lower bound of 2n12n-1 on the rank of an explicit $(\log_{2}(n),n,n)$-tensor (for $n$ a power of two). For border rank the best explicit lower bound is that of [LM19] and is of the form $2.02n$ for an explicit $(n,n,n)$-tensor. To the best of our knowledge, the best tensor rank (or border rank) lower bounds for $(D,n,n)$-tensors with constant $D$ are of the form $3n/2$ [Gri78, Ja’78, Str83, Gri86]. In fact, one can get this bound even with $D=2$ (or $D=3$ for border rank).
@@
To state our main theorem we need one additional notation. Given a set of $n\times n$ matrices ${\cal A}=\{A_{1},\ldots,A_{D}\}$ we define the $(D,n,n)$-tensor $T^{\cal A}$ to be
@@
We prove this theorem in Section 2. An interesting special case of this theorem is that of $(\epsilon n,(1-\epsilon)n)$-dimension spreading maps with $\epsilon$ a small constant. In this case, we can take $D$ to depend only on $\epsilon$ and the rank bound becomes $(2-2\epsilon)n$. We discuss explicit constructions of such dimension-spreading maps in Section 3. We observe that these can be constructed in a black-box manner from dimension-expanders for which there are several known explicit constructions [DS11, DW10, LZ08, BY13, FG15, GRX18]. In particular, using the construction in [BY13] one can get, for any $\epsilon$ and any sufficiently large $n$ an explicit family of $(\epsilon n,(1-\epsilon)n)$-dimension spreading maps over ${\mathbb{F}}^{n}$ of size $D=\textsf{poly}(1/\epsilon)$. [^3] We use the notation $\textsf{poly}(f)=O(f^{C})$ with $C$ some absolute constant. A nice property of this construction is that it is field independent – the same construction works over any field and the defining matrices have only zero and one entries. Hence we get the following corollary.
@@ 3 Constructing dimension-spreading maps
In this section we prove Corollary 1.3 by describing an explicit construction of an $(\epsilon n,(1-\epsilon)n)$-dimension spreading family of constant (depending on $\epsilon$) size. The construction will follow in a black box manner from the related (and well studied) notion of dimension expanders, defined by Barak, Impagliazzo, Shpilka and Wigderson [BISW04] as a linear algebraic analog of expander graphs.
@@
Finally, the most general construction is due to Bourgain and Yehudayoff [BY13] and works over any field. It follows from an explicit construction of a constant degree monotone expander. A degree $D$ monotone graph is a bipartite graph $G$ on vertex set $n]\times[n]$ that can be partitioned into $D$ monotone matchings, which are (partial) matchings in which every pair of edges $(i_{1},j_{1})$ and $(i_{2},j_{2})$ satisfy $i_{1}<i_{2}\,\implies j_{1}<j_{2}$. If $G$ is also a vertex expander (every set $A\subset[n]$ of at most $n/2$ left vertices has at least $(1+\tau)|A|$ neighbors) then $G$ is a degree $D$ monotone expander. It was shown in [DW10] (and implicitly used in [DS11]) that a degree $D$ monotone expander gives a set of $D$ linear maps that are a $\tau$-dimension expander. The linear maps are defined directly from the matchings comprising the graph as follows: Suppose $f:[n]\rightarrow[n]\cup\{\bot\}$ is a partial function representing one of the monotone matchings such that $i_{1}<i_{2}\implies f(i_{1})<f(i_{2})$ (in case both are defined). We define the linear map $A_{f}$ sending $e_{i}$ to $e_{f(i)}$ if $f(i)\neq\bot$ and to 0 otherwise, where $e_{1},\ldots,e_{n}$ is the standard basis of ${\mathbb{F}}^{n}$. It is not hard to show that, over any field, the resulting family of $D$ maps $A_{f}$, taken over all matchings defining $G$, is a dimension-expander whenever the graph $G$ is a monotone expander.
Unlike with ‘standard’ expander graph (without monotonicity), there is no easy probabilistic argument showing existence of constant degree monotone expanders. In [DW10] explicit constructions of monotone expanders with slowly growing (but not constant) $D=D(n)$ were given using an analog of the Zig-Zag product. The construction in [BY13] is the only known construction (or even existence proof) of constant degree monotone expanders and uses a far reaching extension of the Bourgain-Gambourd ‘machine’ [BG07] (used to construct expanding Cayley graphs for finite groups) to the group $SL_{2}({\mathbb{R}})$. Using the dimension expanders of [BY13], which give some (unspecified but explicit) constants $D$ and $\tau$ and using Lemma 3.2 we conclude the proof of Corollary 1.3.

### v2

@@ 1 Overview
An $(n_{1},n_{2},n_{3})$-tensor over $\mathbb{F}$} is a function $T:[n_{1}]\times[n_{2}]\times[n_{3}]\rightarrow{\mathbb{F}}$. We can form two dimensional ‘slices’ of the tensor $T$ by fixing one of the variables. For example, we denote by T(a,,)T(a,*,*) the $n_{2}\times n_{3}$ matrix obtained by fixing $i_{1}$ to $a\in{\mathbb{F}}$. A tensor $T$ has rank one if it is of the form $T(i_{1},i_{2},i_{3})=f_{1}(i_{1})f_{2}(i_{2})f_{3}(i_{3})$ for some functions $f_{j}:[n_{j}]\rightarrow{\mathbb{F}}$. The rank of a general tensor $T$, denoted $R(T)$ is defined to be the minimal $r$ such that $T$ can be written as a sum of $r$ rank-one tensors. [^1] There are other notions of rank for tensors, all generalizing matrix rank, but the one we defined is typically called just ‘tensor rank’. When $\mathbb{F}$} is the real or complex numbers we also define the border rank of $T$, denoted $\underline{R}(T)$ to be the smallest $r$ such that there is a sequence of rank $\leq r$ tensors converging to TT. [^2] One can define border rank also over fields of positive characteristic as being in the Zariski closure of low rank tensors. While it is easy to show the existence of tensors of high rank (as high as $\Omega(n^{2})$ for $(n,n,n)$ tensors) it is a longstanding open problem to construct high-rank tensors explicitly. By explicit construction we mean that, there is a deterministic algorithm that, given $n$, runs in time $\textsf{poly}(n)$ and outputs an $(n,n,n)$-tensor $T_{n}$ of the required rank (more formally, an explicit construction produces an infinite family of tensors, not a single one). The difficulty in constructing explicit tensors of high rank stems from the fact that such constructions would lead to lower bounds on arithmetic circuits, a notoriously difficult open problem in theoretical computer science (see [SY-survey] for more details on this connection).
The currently best known explicit lower bounds on tensor rank are those in [AFT11] and are of the form $3n-O(\log_{2}(n))$ for an $(n,n,n)$-tensor (over any field $\mathbb{F}$}). Their work also includes a lower bound of 2n12n-1 on the rank of an explicit $(\log_{2}(n),n,n)$-tensor (for $n$ a power of two). It is not stated in [AFT11], but one can also slightly modify their Theorem 5.6 to construct a $(C,n,n)$-tensor with rank $(2-2^{-c})n$. As far as we can tell, the above mentioned results from [AFT11] do not extend to border rank.
For border rank the best explicit lower bound is that of [LM19] and is of the form $2.02n$ for an explicit $(n,n,n)$-tensor. To the best of our knowledge, the best border rank lower bounds for $(D,n,n)$-tensors with constant $D$ are of the form $3n/2$ [Gri78, JaJa78, Str83, Grie86]. In fact, one can get this bound even with $D=3$.
@@
To state our main theorem, we need one additional notation. Given a set of $n\times n$ matrices ${\cal A}=\{A_{1},\ldots,A_{D}\}$ we define the $(D,n,n)$-tensor $T^{\cal A}$ to be
@@
We prove this theorem in Section 2. An interesting special case of this theorem is that of $(\epsilon n,(1-\epsilon)n)$-dimension spreading maps with $\epsilon$ a small constant. In this case, we can take $D$ to depend only on $\epsilon$ and the rank bound becomes $(2-2\epsilon)n$. We discuss explicit constructions of such dimension-spreading maps in Section 3. We observe that these can be constructed in a black-box manner from dimension-expanders for which there are several known explicit constructions [DS07, DW10, LZ08, BY13, FG15, GRX18]. In particular, using the construction in [BY13] one can get, for any $\epsilon$ and any sufficiently large $n$ an explicit family of $(\epsilon n,(1-\epsilon)n)$-dimension spreading maps over ${\mathbb{F}}^{n}$ of size $D=\textsf{poly}(1/\epsilon)$. [^3] We use the notation $\textsf{poly}(f)=O(f^{C})$ with $C$ some absolute constant. A nice property of this construction is that it is field independent – the same construction works over any field and the defining matrices have only zero and one entries. Hence, we get the following corollary.
@@ 3 Constructing dimension-spreading maps
In this section we prove Corollary 1.3 by describing an explicit construction of an $(\epsilon n,(1-\epsilon)n)$-dimension spreading family of constant (depending on $\epsilon$) size. The construction will follow in a black box manner from the related (and well studied) notion of dimension expanders, defined by Barak, Impagliazzo, Shpilka and Wigderson [BISW] as a linear algebraic analog of expander graphs.
@@
Finally, the most general construction is due to Bourgain and Yehudayoff [BY13] and works over any field. It follows from an explicit construction of a constant degree monotone expander. A degree $D$ monotone graph is a bipartite graph $G$ on vertex set $n]\times[n]$ that can be partitioned into $D$ monotone matchings, which are (partial) matchings in which every pair of edges $(i_{1},j_{1})$ and $(i_{2},j_{2})$ satisfy $i_{1}<i_{2}\,\implies j_{1}<j_{2}$. If $G$ is also a vertex expander (every set $A\subset[n]$ of at most $n/2$ left vertices has at least $(1+\tau)|A|$ neighbors) then $G$ is a degree $D$ monotone expander. It was shown in [DW10] (and implicitly used in [DS07]) that a degree $D$ monotone expander gives a set of $D$ linear maps that are a $\tau$-dimension expander. The linear maps are defined directly from the matchings comprising the graph as follows: Suppose $f:[n]\rightarrow[n]\cup\{\bot\}$ is a partial function representing one of the monotone matchings such that $i_{1}<i_{2}\implies f(i_{1})<f(i_{2})$ (in case both are defined). We define the linear map $A_{f}$ sending $e_{i}$ to $e_{f(i)}$ if $f(i)\neq\bot$ and to 0 otherwise, where $e_{1},\ldots,e_{n}$ is the standard basis of ${\mathbb{F}}^{n}$. It is not hard to show that, over any field, the resulting family of $D$ maps $A_{f}$, taken over all matchings defining $G$, is a dimension-expander whenever the graph $G$ is a monotone expander.
Unlike with ‘standard’ expander graph (without monotonicity), there is no easy probabilistic argument showing the existence of constant degree monotone expanders. In [DW10] explicit constructions of monotone expanders with slowly growing (but not constant) $D=D(n)$ were given using an analog of the Zig-Zag product. The construction in [BY13] is the only known construction (or even existence proof) of constant degree monotone expanders and uses a far reaching extension of the Bourgain-Gambourd ‘machine’ [BG07] (used to construct expanding Cayley graphs for finite groups) to the group $SL_{2}({\mathbb{R}})$. Using the dimension expanders of [BY13], which give some (unspecified but explicit) constants $D$ and $\tau$ and using Lemma 3.2 we conclude the proof of Corollary 1.3.